# Global well-posedness for small data in a 3D temperature-velocity model with Dirichlet boundary noise

### v2

@@ 3.2. The remainder temperature equation
For any $0<\delta_{\theta}\leq\overline{\delta}_{\theta}$, any velocity field
@@
To simplify the notation, we set $\varepsilon=1$ and drop the dependence on $\varepsilon$ of $\xi^{\varepsilon}$. We start by recalling that from [6, Theorem 7.4], given $s\geq 0$, the pointwise multiplication
@@
    $s_{1}+s_{2}>s+3\left(\frac{1}{p_{1}}+\frac{1}{p_{2}}-\frac{1}{p}\right).$
@@
By the norm equivalence between $W^{\frac{1}{2}+\lambda,2}(\mathcal{D})$ and $H^{\frac{1}{2}+\lambda}(\mathcal{D})$, and by Corollary 3.1, we have $u_{t}\in W^{\frac{1}{2}+\lambda,2}\left(\mathcal{D}\right)$ a.e., and since $\varphi\in L^{6}\left(\mathcal{D}\right)$, we deduce $\nabla(-\Delta)^{-\gamma-\frac{1}{2}}\varphi\in W^{2\gamma,6}\left(\mathcal{D}\right)$. Hence, (3.15) follows from the pointwise multiplication embedding
@@
Let $s\in[0,\frac{1}{2})$, $p>1$ and $\gamma>\frac{1}{4}$ be given. Assume that
@@
Fix $0<\delta_{\theta}\leq\overline{\delta}_{\theta}=\min\{\lambda,2\gamma-\frac{1}{2}\},$ and consider $\alpha_{\theta}:=\frac{1}{4}+\frac{\delta_{\theta}}{2}.$ Define
@@
where $C=C(s,p,\gamma,\lambda,\delta_{\theta},T)>0$. Lastly, given $s\in[0,\frac{1}{2}),p>1$ as before, there always exists $\gamma=\gamma(s)>\frac{1}{4}$ and $\lambda=\lambda(s,p)\in(0,1)$ such that (3.17) and (3.18) hold.
@@
and the right-hand side is integrable on $(0,r)$ by the same computation as in (3.21). Hence, by the dominated convergence theorem, $\|I_{2}(t,r)\|_{L^{\frac{6}{5}}}\to 0$ as $t\downarrow r$. We deduce that $\|\rho_{t}-\rho_{r}\|_{L^{\frac{6}{5}}(\mathcal{D})}\to 0$ for every $r\in(0,T]$, that is $\rho\in C(0,T;L^{\frac{6}{5}}(\mathcal{D}))$. Consequently, $\chi_{t}=(-\Delta)^{-\frac{s}{2}}\rho_{t}$ belongs to $C(0,T;W^{s,\frac{6}{5}}(\mathcal{D})\bigr)$.
Step 4. Existence of $\gamma,\lambda$, given $s,p$. Assume that, as before, $s\in[0,\frac{1}{2})$ and $p>1$ are given. We now check that we can always find $\gamma=\gamma(s)>\frac{1}{4}$ and $\lambda=\lambda(s,p)$ such that (3.17) and (3.18) are satisfied, and thus this corollary is non-trivial. Indeed, with the notation of Step $1$, from $a(\frac{1}{4})=\frac{3}{4}+\frac{s}{2}<1$, it follows that there exists $\gamma=\gamma(s)>\frac{1}{4}$ sufficiently close to $\frac{1}{4}$ such that $a(s,\gamma)<1$, that is (3.17). Hence, also p(1a)>0p(1-a)>0 holds, and we may choose $\lambda=\lambda(s,p)\in(0,1)$ so small that $\lambda<p(1-a),$ that is (3.18). ∎
@@
From the proof of Corollary 3.3, it follows that if $s\in[0,\frac{1}{2}),p>1$, $\gamma>\frac{1}{4}$ and $\lambda\in(0,1)$ satisfy (3.17) and (3.18), then
@@
For simplicity we set $\varepsilon=1$ and drop the superscript, writing $\zeta$ and $Z$ instead of $\zeta^{\varepsilon}$ and $Z^{\varepsilon}$. We also use the notation of Corollary 3.3 and Proposition 2.1. Let $s\in[0,\frac{1}{2})$ and $p>2$ be given.
Step $1$. Choices of $\gamma,q,q^{\prime},\lambda,\overline{\delta}_{\theta},\overline{\delta}_{u}.$
Choice of $\gamma=\gamma(s,p)>\frac{1}{4}.$ We choose $\gamma=\frac{1}{4}+\varepsilon_{\gamma}$, with $\varepsilon_{\gamma}>0$ sufficiently small such that
(3.22) $\left\{\begin{aligned} a(s,\gamma):=\frac{1}{2}+\gamma+\frac{s}{2}&<1,\\ \lambda_{\max}(s,p,\gamma):=p(1-a(s,\gamma))&>\frac{1}{2}-s.\end{aligned}\right$.
The condition (3.22)1 is satisfied, up to choosing $\varepsilon_{\gamma}$ sufficiently small, since
$\lim\limits_{\gamma\downarrow\frac{1}{4}}a(s,\gamma)=\frac{3}{4}+\frac{s}{2}<1,$
where in the last inequality we use $s<\frac{1}{2}$. For the condition (3.22)2, note that if $\gamma\in(\frac{1}{4},\frac{1-s}{2})$, then $0<\frac{1}{2}-\gamma-\frac{s}{2}=1-a(s,\gamma)$. Further, since $p>2$ by hypothesis, we have
$\lim\limits_{\gamma\downarrow\frac{1}{4}}\lambda_{\max}(s,p,\gamma)=\frac{p}{2}(\frac{1}{2}-s)>\frac{1}{2}-s.$
Choice of $q=q(s)$ and $q^{\prime}=q^{\prime}(s)$. Let $q=q(s):=\frac{6}{5-2s}.$ With this choice, the Sobolev embedding
(3.23) $W^{s,\frac{6}{5}}(\mathcal{D})\hookrightarrow L^{q(s)}(\mathcal{D})$
holds, a condition that we will later need in the proof. Further, denoting by $q^{\prime}$ the conjugate exponent of $q$, it holds
(3.24) $1-\frac{3}{q^{\prime}}<\lambda_{\max}(s,\gamma,p).$
Indeed
$1-\frac{3}{q^{\prime}(s)}=-2+\frac{3}{q(s)}=\frac{1}{2}-s,$
and (3.24) then follows from (3.22)2.
Choice of $\lambda=\lambda(s).$ Define
$\lambda_{\min}(s):=\frac{1}{2}-s=1-\frac{3}{q^{\prime}(s)}.$
and note that $\lambda_{\min}\in(0,1)$ since $s\in[0,\frac{1}{2})$. Further, $\lambda_{\min}<\lambda_{\max}$. Thus, we can choose $\lambda$ sufficiently close to $\lambda_{\min}$ such that
$\lambda=\lambda(s)\in(\lambda_{\min},\lambda_{\max})\cap(0,1).$
With this choice of $\lambda$, we claim that the pointwise multiplication embedding
(3.25) $W^{\frac{1}{2}+\lambda(s),2}(\mathcal{D})\times W^{2\gamma(s),6}(\mathcal{D})\hookrightarrow L^{q^{\prime}(s)}(\mathcal{D})$
holds, another condition that we will need later in the proof. Indeed, thanks to [6, Theorem 7.4], it is sufficient to check that
  1. (i)
    $\frac{1}{2}+\lambda\geq 0$, $2\gamma\geq 0$;
  2. (ii)
    $\frac{1}{2}+\lambda-\frac{3}{2}\geq-\frac{3}{q^{\prime}}$, $2\gamma-\frac{1}{2}\geq-\frac{3}{q^{\prime}}$;
  3. (iii)
    $\frac{1}{2}+\lambda+2\gamma-3(\frac{1}{2}+\frac{1}{6})>-\frac{3}{q^{\prime}}.$
The previous conditions are equivalent to
  1. (a)
    $\lambda\geq-\frac{1}{2},$ $\gamma\geq 0$;
  2. (b)
    $\lambda\geq 1-\frac{3}{q^{\prime}},$ $2\gamma\geq\frac{1}{2}-\frac{3}{q^{\prime}};$
  3. (c)
    $\lambda+2\gamma\geq\frac{3}{2}-\frac{3}{q^{\prime}}$.
The conditions in (a) and (b) follow from the choice of $\lambda,\gamma$, in particular thanks to $\lambda>\lambda_{\min}>0$ and $\gamma>\frac{1}{4}.$ Lastly, since $\lambda+2\gamma>\lambda+\frac{1}{2}$, then (c) is satisfied if $\lambda+\frac{1}{2}>\frac{3}{2}-\frac{3}{q^{\prime}}$, which is true thanks to the choice $\lambda>\lambda_{\min}.$
Choice of $\overline{\delta}_{\theta}$ and $\overline{\delta}_{u}$. With $\gamma$ and $\lambda$ fixed as above, we define the thresholds
@@
$\overline{\delta}_{\theta}(s,p):=\min\left\{\lambda,2\gamma-\frac{1}{2}\right\},\quad\overline{\delta}_{u}(s,p):=\min\left\{\lambda_{\max}(s,p,\gamma)-\lambda,1-\lambda\right\},$
and note that $\overline{\delta}_{\theta}$ and $\overline{\delta}_{u}$ are the same as in Corollary 3.3. Further, both are positive by the choices made above. We assume $0<\delta_{\theta}\leq\overline{\delta}_{\theta}$ and $0<\delta_{u}<\overline{\delta}_{u}$. Since $\delta_{\theta}\leq\overline{\delta}_{\theta}$ and $\delta_{u}<\overline{\delta}_{u}$, the assumptions of Corollary 3.3 are satisfied. Furthermore, from Remark 3.4, it follows that
(3.26) $(\gamma+\frac{s}{2}+\frac{1}{2})\frac{1}{1-\frac{\lambda}{p}-\frac{\delta_{u}}{p}}<1.$
@@
$(-\Delta)^{-\gamma-\frac{1}{2}}\mathrm{div}\left(u_{t}\zeta_{t}\right)\in L^{\frac{p}{\lambda+\delta_{u}}}(0,T_{1};L^{\frac{6}{5}}(\mathcal{D})).$
As in the proof of Lemma 3.2, given $\varphi\in L^{6}(\mathcal{D})$ we have
$\left\langle(-\Delta)^{-\gamma-\frac{1}{2}}\mathrm{div}\left(u_{t}\zeta_{t}\right),\varphi\right\rangle$ $=\langle\mathrm{div}(u_{t}\zeta_{t}),(-\Delta)^{-\gamma-\frac{1}{2}}\varphi\rangle$
$=-\langle u_{t}\zeta_{t},\nabla(-\Delta)^{-\gamma-\frac{1}{2}}\varphi\rangle$
$=-\langle\zeta_{t},u_{t}\cdot\nabla(-\Delta)^{-\gamma-\frac{1}{2}}\varphi\rangle.$
We now consider the conjugate exponents $q=q(s),q^{\prime}=q^{\prime}(s)\in(1,+\infty)$ chosen in Step 1. We know that $u\in L^{\frac{p}{\lambda+\delta_{u}}}\left(0,T;W^{\frac{1}{2}+\lambda,2}(\mathcal{D})\right)$ thanks to the norm equivalence between $W^{\frac{1}{2}+\lambda,2}(\mathcal{D})$ and $H^{\frac{1}{2}+\lambda}(\mathcal{D})$, and Corollary 3.1 (note that its hypotheses are satisfied since $\delta_{u}\in(0,1)$, $\lambda\in(0,1-\delta_{u})$). Since the pointwise multiplication embedding
(3.28) $\begin{split}W^{\frac{1}{2}+\lambda(s),2}(\mathcal{D})\times W^{2\gamma(s),6}(\mathcal{D})&\hookrightarrow L^{q^{\prime}(s)}(\mathcal{D})\\ (u_{t},\nabla(-\Delta)^{-\gamma-\frac{1}{2}}\varphi)&\mapsto u_{t}\cdot\nabla(-\Delta)^{-\gamma-\frac{1}{2}}\varphi\end{split}$
holds thanks to the condition verified after (3.25), we get that there exists $C=C(\lambda(s),\gamma(s))=C(s)>0$ such that
$\left\|u_{t}\cdot\nabla(-\Delta)^{-\gamma-\frac{1}{2}}\varphi\right\|_{L^{q^{\prime}}(\mathcal{D})}\leq C\left\|u_{t}\right\|_{W^{\frac{1}{2}+\lambda,2}(\mathcal{D})}\left\|\nabla(-\Delta)^{-\gamma-\frac{1}{2}}\varphi\right\|_{W^{2\gamma,6}(\mathcal{D})},$
and thus
$\left\|(-\Delta)^{-\gamma-\frac{1}{2}}\mathrm{div}\left(u_{t}\zeta_{t}\right)\right\|_{L^{\frac{6}{5}}(\mathcal{D})}$ $=\sup_{\|\varphi\|_{6}\leq 1}\left\langle(-\Delta)^{-\gamma-\frac{1}{2}}\mathrm{div}\left(u_{t}\zeta_{t}\right),\varphi\right\rangle$
$\leq C\left\|\zeta_{t}\right\|_{L^{q}(\mathcal{D})}\left\|u_{t}\right\|_{W^{\frac{1}{2}+\lambda,2}(\mathcal{D})}\sup_{\|\varphi\|_{6}\leq 1}\left\|\nabla(-\Delta)^{-\gamma-\frac{1}{2}}\varphi\right\|_{W^{2\gamma,6}(\mathcal{D})}$
$\leq C\left\|\zeta_{t}\right\|_{L^{q}(\mathcal{D})}\left\|u_{t}\right\|_{W^{\frac{1}{2}+\lambda,2}(\mathcal{D})}\sup_{\|\varphi\|_{6}\leq 1}\left\|\varphi\right\|_{L^{6}(\mathcal{D})}$
$=C\left\|\zeta_{t}\right\|_{L^{q}(\mathcal{D})}\left\|u_{t}\right\|_{W^{\frac{1}{2}+\lambda,2}(\mathcal{D})}.$
Further, since $\zeta\in X=C(0,T_{1};W^{s,\frac{6}{5}}(\mathcal{D}))\hookrightarrow C(0,T_{1};L^{q}(\mathcal{D}))$ thanks to (3.23), we can write
(3.29) $\int_{0}^{T_{1}}\left\|(-\Delta)^{-\gamma-\frac{1}{2}}\mathrm{div}\left(u_{t}\zeta_{t}\right)\right\|_{L^{\frac{6}{5}}(\mathcal{D})}^{\frac{p}{\lambda+\delta_{u}}}dt\leq C\left(\sup_{t\in[0,T_{1}]}\left\|\zeta_{t}\right\|_{L^{q}(\mathcal{D})}\right)^{\frac{p}{\lambda+\delta_{u}}}\int_{0}^{T_{1}}\left\|u_{t}\right\|_{W^{\frac{1}{2}+\lambda,2}(\mathcal{D})}^{\frac{p}{\lambda+\delta_{u}}}dt<\infty.$
The thesis of this step follows at this point since $u\in L^{\frac{p}{\lambda+\delta_{u}}}(0,T;W^{\frac{1}{2}+\lambda,2}(\mathcal{D}))$.
@@
Step 5. We conclude the proof by showing that $\Psi$ is a contraction from $X$ to itself, up to choosing $T_{1}$ small enough. Consider $\zeta_{1},\zeta_{2}\in X$ and their difference $\zeta:=\zeta_{1}-\zeta_{2}$, then, from

### v3

@@ 3.2. The remainder temperature equation
For any $0<\delta_{\theta}<\overline{\delta}_{\theta}$, any velocity field
@@
To simplify the notation, we set $\varepsilon=1$ and drop the dependence on $\varepsilon$ of $\xi^{\varepsilon}$. We start by recalling that, by [6, Theorem A.1] (which we can apply after extending functions on $\mathcal{D}$ to $\mathbb{R}^{3}$ via a bounded extension operator), the pointwise multiplication
@@
    $s_{1}+s_{2}>s+3\left(\frac{1}{p_{1}}+\frac{1}{p_{2}}-\frac{1}{p}\right)$;
  4. (iv)
    if $s_{i}=s\notin\mathbb{N}$ for some $i\in\{1,2\}$, then one must additionally assume $p_{i}\leq p$.
@@
By the norm equivalence between $W^{\frac{1}{2}+\lambda,2}(\mathcal{D})$ and $H^{\frac{1}{2}+\lambda}(\mathcal{D})$, and by Corollary 3.1, we have $u_{t}\in W^{\frac{1}{2}+\lambda,2}\left(\mathcal{D}\right)$ a.e., and since $\varphi\in L^{6}\left(\mathcal{D}\right)$, we deduce $\nabla(-\Delta)^{-\gamma-\frac{1}{2}}\varphi\in W^{2\gamma,6}\left(\mathcal{D}\right)$. Thus (3.14) follows from the pointwise multiplication embedding
@@
Since in general $s=\frac{1}{2}+\delta_{\theta}\notin\mathbb{N}$ and here $p_{2}=6>p=2$, condition (iv) in [6, Theorem A.1] forbids the borderline case $s_{2}=s$. Hence we impose
$s_{2}>s\quad\Longleftrightarrow\quad 2\gamma>\frac{1}{2}+\delta_{\theta}\quad\Longleftrightarrow\quad\delta_{\theta}<2\gamma-\frac{1}{2}.$
In contrast, the endpoint $\delta_{\theta}=\lambda$ is admissible since then $s_{1}=s$ but $p_{1}=p=2$.
@@
Let $s\in(0,\frac{1}{2})$, $p>1$ and $\gamma>\frac{1}{4}$ be given. Assume that
@@
Fix $0<\delta_{\theta}<\overline{\delta}_{\theta}=\min\{\lambda,2\gamma-\frac{1}{2}\},$ and consider $\alpha_{\theta}:=\frac{1}{4}+\frac{\delta_{\theta}}{2}.$ Define
@@
where $C=C(s,p,\gamma,\lambda,\delta_{\theta},T)>0$. Lastly, given $s\in(0,\frac{1}{2}),p>1$ as before, there always exists $\gamma=\gamma(s)>\frac{1}{4}$ and $\lambda=\lambda(s,p)\in(0,1)$ such that (3.16) and (3.17) hold.
@@
and the right-hand side is integrable on $(0,r)$ by the same computation as in (3.20). Hence, by the dominated convergence theorem, $\|I_{2}(t,r)\|_{L^{\frac{6}{5}}}\to 0$ as $t\downarrow r$. We deduce that $\|\rho_{t}-\rho_{r}\|_{L^{\frac{6}{5}}(\mathcal{D})}\to 0$ for every $r\in(0,T]$, that is $\rho\in C(0,T;L^{\frac{6}{5}}(\mathcal{D}))$. Consequently, $\chi_{t}=(-\Delta)^{-\frac{s}{2}}\rho_{t}$ belongs to $C(0,T;W^{s,\frac{6}{5}}(\mathcal{D}))$.
Step 4. Existence of $\gamma,\lambda$, given $s,p$. Assume that, as before, $s\in(0,\frac{1}{2})$ and $p>1$ are given. We now check that we can always find $\gamma=\gamma(s)>\frac{1}{4}$ and $\lambda=\lambda(s,p)$ such that (3.16) and (3.17) are satisfied, and thus this corollary is non-trivial. Indeed, with the notation of Step $1$, from $a(\frac{1}{4})=\frac{3}{4}+\frac{s}{2}<1$, it follows that there exists $\gamma=\gamma(s)>\frac{1}{4}$ sufficiently close to $\frac{1}{4}$ such that $a(s,\gamma)<1$, that is (3.16). Hence, also p(1a)>0p(1-a)>0 holds, and we may choose $\lambda=\lambda(s,p)\in(0,1)$ so small that $\lambda<p(1-a),$ that is (3.17). ∎
@@
From the proof of Corollary 3.3, it follows that if $s\in(0,\frac{1}{2}),p>1$, $\gamma>\frac{1}{4}$ and $\lambda\in(0,1)$ satisfy (3.16) and (3.17), then
@@
For simplicity we set $\varepsilon=1$ and drop the superscript, writing $\zeta$ and $Z$ instead of $\zeta^{\varepsilon}$ and $Z^{\varepsilon}$. Let $s\in(0,\frac{1}{2})$ and $p>4$ be given.
Step 1. Setup of parameters and auxiliary embeddings. We work with the parameters $\gamma,\lambda,\overline{\delta}_{\theta},\overline{\delta}_{u}$ fixed in the statement of the theorem. In particular,
$a(s,\gamma):=\frac{1}{2}+\gamma+\frac{s}{2}<1,\qquad\lambda_{\max}:=p(1-a(s,\gamma)),$
and, by (2.9), we have $\lambda_{\max}>\frac{1}{2}-s$. Moreover, $\lambda_{\max}-(1-2s)>0$ by the choice of $\varepsilon_{\gamma}$ in (2.8) as shown in (2.10). We have set
$\lambda:=(\frac{1}{2}-s)+\varepsilon_{\lambda},\qquad 0<\varepsilon_{\lambda}<\min\{2s,\lambda_{\max}-(1-2s)\}.$
Note that
$\lambda<(\tfrac{1}{2}-s)+2s=\tfrac{1}{2}+s<1,\qquad\lambda<(\tfrac{1}{2}-s)+\lambda_{\max}-(1-2s)=\lambda_{\max}-(\tfrac{1}{2}-s)<\lambda_{\max}.$
Hence $\lambda\in(\tfrac{1}{2}-s,\min\{1,\lambda_{\max}\})$. The thresholds are
@@
$\overline{\delta}_{\theta}=\min\left\{\lambda,2\gamma-\frac{1}{2}\right\},\qquad\overline{\delta}_{u}=\min\left\{\lambda_{\max}-\lambda,1-\lambda\right\}.$
Finally, fix $0<\delta_{\theta}<\overline{\delta}_{\theta}$ and $0<\delta_{u}<\overline{\delta}_{u}$.
Choice of $q,q^{\prime}$. Set
$q:=\frac{6}{5-2s},\qquad\frac{1}{q}+\frac{1}{q^{\prime}}=1.$
Then the Sobolev embedding
(3.21) $W^{s,\frac{6}{5}}(\mathcal{D})\hookrightarrow L^{q}(\mathcal{D})$
holds, and a direct computation gives
(3.22) $1-\frac{3}{q^{\prime}}=\frac{1}{2}-s.$
Product embedding. Since $\gamma>\frac{1}{4}$, the Sobolev embedding yields
(3.23) $W^{2\gamma,6}(\mathcal{D})\hookrightarrow L^{\infty}(\mathcal{D}).$
Moreover, again by the Sobolev embedding,
(3.24) $W^{\frac{1}{2}+\lambda,2}(\mathcal{D})\hookrightarrow L^{r_{\lambda}}(\mathcal{D}),\qquad r_{\lambda}:=\frac{3}{1-\lambda}.$
Recalling that $q^{\prime}=\tfrac{6}{1+2s},$ it can be checked that $r_{\lambda}\geq q^{\prime}$ is equivalent to $\lambda\geq\tfrac{1}{2}-s$. Hence, by (3.24) and the embedding $L^{r_{\lambda}}(\mathcal{D})\hookrightarrow L^{q^{\prime}}(\mathcal{D})$ on bounded $\mathcal{D}$,
(3.25) $W^{\frac{1}{2}+\lambda,2}(\mathcal{D})\hookrightarrow L^{q^{\prime}}(\mathcal{D}).$
Combining (3.23), (3.25) with Hölder’s inequality yields the continuous multiplication map
(3.26) $W^{\frac{1}{2}+\lambda,2}(\mathcal{D})\times W^{2\gamma,6}(\mathcal{D})\hookrightarrow L^{q^{\prime}}(\mathcal{D}).$
In particular, for every $\varphi\in L^{6}(\mathcal{D})$,
(3.27) $\bigl\|u_{t}\cdot\nabla(-\Delta)^{-\gamma-\frac{1}{2}}\varphi\bigr\|_{L^{q^{\prime}}(\mathcal{D})}\leq C\,\|u_{t}\|_{W^{\frac{1}{2}+\lambda,2}(\mathcal{D})}\,\|\varphi\|_{L^{6}(\mathcal{D})}.$
Time integrability condition. Since $\delta_{u}<\overline{\delta}_{u}\leq 1-\lambda$, we have $\lambda\leq 1-\delta_{u}$ and therefore Corollary 3.1 implies
(3.28) $u\in L^{\frac{p}{\lambda+\delta_{u}}}\bigl(0,T;W^{\frac{1}{2}+\lambda,2}(\mathcal{D})\bigr).$
Moreover, $\delta_{u}<\overline{\delta}_{u}\leq\lambda_{\max}-\lambda$ gives $\lambda+\delta_{u}<\lambda_{\max}=p(1-a(s,\gamma))$, and hence, by Remark 3.4,
(3.29) $\Bigl(\gamma+\frac{s}{2}+\frac{1}{2}\Bigr)\frac{1}{1-\frac{\lambda}{p}-\frac{\delta_{u}}{p}}<1.$
@@
(3.31) $(-\Delta)^{-\gamma-\frac{1}{2}}\mathrm{div}\left(u_{t}\zeta_{t}\right)\in L^{\frac{p}{\lambda+\delta_{u}}}\bigl(0,T_{1};L^{\frac{6}{5}}(\mathcal{D})\bigr).$
Let $\varphi\in L^{6}(\mathcal{D})$. As in the proof of Lemma 3.2, using integration by parts and the self-adjointness of $\Delta$, we obtain
$\left\langle(-\Delta)^{-\gamma-\frac{1}{2}}\mathrm{div}\left(u_{t}\zeta_{t}\right),\varphi\right\rangle$ $=\left\langle\mathrm{div}(u_{t}\zeta_{t}),(-\Delta)^{-\gamma-\frac{1}{2}}\varphi\right\rangle$
$=-\left\langle u_{t}\zeta_{t},\nabla(-\Delta)^{-\gamma-\frac{1}{2}}\varphi\right\rangle=-\left\langle\zeta_{t},u_{t}\cdot\nabla(-\Delta)^{-\gamma-\frac{1}{2}}\varphi\right\rangle.$
Let $q,q^{\prime}$ be the conjugate exponents introduced in Step 1, namely $q=\frac{6}{5-2s}$ and $1/q+1/q^{\prime}=1$. By Hölder’s inequality,
$\bigl|\left\langle\zeta_{t},u_{t}\cdot\nabla(-\Delta)^{-\gamma-\frac{1}{2}}\varphi\right\rangle\bigr|\leq\|\zeta_{t}\|_{L^{q}(\mathcal{D})}\;\|u_{t}\cdot\nabla(-\Delta)^{-\gamma-\frac{1}{2}}\varphi\|_{L^{q^{\prime}}(\mathcal{D})}.$
Using the product estimate (3.27) from Step 1, we infer
$\|u_{t}\cdot\nabla(-\Delta)^{-\gamma-\frac{1}{2}}\varphi\|_{L^{q^{\prime}}(\mathcal{D})}\leq C\,\|u_{t}\|_{W^{\frac{1}{2}+\lambda,2}(\mathcal{D})}\,\|\varphi\|_{L^{6}(\mathcal{D})},$
and therefore
$\bigl|\left\langle(-\Delta)^{-\gamma-\frac{1}{2}}\mathrm{div}\left(u_{t}\zeta_{t}\right),\varphi\right\rangle\bigr|\leq C\,\|\zeta_{t}\|_{L^{q}(\mathcal{D})}\;\|u_{t}\|_{W^{\frac{1}{2}+\lambda,2}(\mathcal{D})}\;\|\varphi\|_{L^{6}(\mathcal{D})}.$
Taking the supremum over $\|\varphi\|_{L^{6}}\leq 1$ yields
(3.32) $\|(-\Delta)^{-\gamma-\frac{1}{2}}\mathrm{div}(u_{t}\zeta_{t})\|_{L^{\frac{6}{5}}(\mathcal{D})}\leq C\,\|\zeta_{t}\|_{L^{q}(\mathcal{D})}\;\|u_{t}\|_{W^{\frac{1}{2}+\lambda,2}(\mathcal{D})}.$
Since $\zeta\in X=C(0,T_{1};W^{s,\frac{6}{5}}(\mathcal{D}))$, the Sobolev embedding (3.21) implies $\sup_{t\in[0,T_{1}]}\|\zeta_{t}\|_{L^{q}}<\infty$. Hence, raising (3.32) to the power $\frac{p}{\lambda+\delta_{u}}$ and integrating, we obtain
(3.33) $\int_{0}^{T_{1}}\|(-\Delta)^{-\gamma-\frac{1}{2}}\mathrm{div}(u_{t}\zeta_{t})\|_{L^{\frac{6}{5}}}^{\frac{p}{\lambda+\delta_{u}}}\,dt\leq C\Bigl(\sup_{t\in[0,T_{1}]}\|\zeta_{t}\|_{L^{q}}\Bigr)^{\frac{p}{\lambda+\delta_{u}}}\int_{0}^{T_{1}}\|u_{t}\|_{W^{\frac{1}{2}+\lambda,2}}^{\frac{p}{\lambda+\delta_{u}}}\,dt.$
The last integral is finite by (3.28), i.e. $u\in L^{\frac{p}{\lambda+\delta_{u}}}(0,T_{1};W^{\frac{1}{2}+\lambda,2}(\mathcal{D}))$. This proves (3.31)
@@
Step 5. We conclude the proof by showing that $\Psi$ is a contraction in $X=C(0,T_{1};W^{s,\frac{6}{5}}(\mathcal{D}))$ for $T_{1}>0$ sufficiently small. Let $\zeta_{1},\zeta_{2}\in X$ and set $\zeta:=\zeta_{1}-\zeta_{2}$. From